Draw the graph of the function f, defined by f(x)= |x- 6| + |5 - x| ; x∈ [2,8]

1
Expert's answer
2022-06-08T11:24:56-0400

"f(x)=|x-6|+|5-x|"

"|x-6|=(x-6), if (x-6)\geqslant0\iff x\geqslant6"

"|x-6|=-(x-6), if (x-6)<0\iff x<6"

"|5-x|=(5-x), if (5-x)\geqslant0\iff x\leqslant5"

"|5-x|=-(5-x), if (5-x)<0\u27fax>5"

1) "x\in[2,5]"

"f(x)=-(x-6)+(5-x)=-2x+11"

2) "x\in(5, 6)"

"f(x)=-(x-6)-(5-x)=-x+6-5+x=1"

3) "x\in[6, 8]"

"f(x)=(x-6)-(5-x)=2x-11"

"f(x)=(x-6)-(5-x)=2x-11"